CALCULATE THE MOLARITY OF 30 GM OF CO[NO3]2.6H2O IN 4.3 L OF SOLUTION

Grade:

2 Answers

Gurwinder Kaur
13 years ago

Mass of CO[NO3]2.6H2O = 30gm

Volume of solution = 4.3L

Molar mass of CO[NO3]2.6H2O = 291 g/mol

no. of moles of solute = 30/291 mol

Molarity of solution = no. of moles of CO[NO3]2.6H2O/ Volume of sol. in litres

                                     = 30/(291*4.3)

                                     = 0.0239 molar

plz approve if it is helpful...........

Joseph Ogar
6 years ago
n = CV
n = no. of moles = mass/molar mass
Therefore n = 30/291 
V = Volume in L = 4.3 L
C = Concentration in mol dm-3 or Molarity 
    = n/V = (30/291)/4.3
    = 0.0239 M